OpenStudy (anonymous):

Part 1: Find the solution to the equation 64^(3 + x) = 4^(2x)? Please explain!!

OpenStudy (anonymous):

\[64^{(3 +x)}=4^{2x}\]

Parth (parthkohli):

We'll get the same, and we know that 64 can be expressed as 4 cubed. \(\Large \color{purple}{\rightarrow (4^3)^{3 + x} = 4^{2x} }\) Multiply the exponents \(\Large \color{purple}{\rightarrow 4^{9 + 3x} = 4^{2x} }\) We can make a normal equation \(\Large \color{purple}{\rightarrow 9 + 3x = 2x }\) \(\Large \color{purple}{\rightarrow 9 = -x }\) \(\Large \color{purple}{\rightarrow x = -9 }\)